# Recursive Inference Scaling: A Winning Path to Scalable Inference in Language and Multimodal Systems

### v2

@@ 3 Experimental Results
In this section, we study the impact of various parameter-sharing strategies in language modeling, following the taxonomy introduced in Section 2. We show how RINS emerges as a clear winner. All experiments are carried out using the Big Vision codebase (Beyer et al., 2022).
Metric Architecture
BL RAO RINS
Winogrande 53.3% 53.0% 53.3%
OpenBookQA 37.7% 38.3% 38.9%
BoolQ 47.7% 41.6% 52.5%
PIQA 54.5% 68.6% 69.5%
SIQA 39.4% 40.2% 41.3%
Table 1: Zero-shot evaluation in downstream common sense reasoning tasks. All models are 600M parameters, pretrained on the compute-equivalent of 500B tokens in the baseline (BL). The best signatures for RAO and RINS in Figure 4 are used in this evaluation. See Appendix C for details on the evaluation setup.
@@
First, we train a decoder-only transformer language model (Vaswani et al., 2017) with relative position embeddings and sequence packing. We use C4/English tokenizer with a vocabulary size of 32K. The model is trained on a mixture of C4 (Raffel et al., 2020) and SlimPajama (Soboleva et al., 2023) with equal weight using a batch size 1,024 and context length 1,024. The non-recursive baseline is trained for 200K steps, which amounts to 200B training tokens. Other recursive models are trained on fewer tokens in order to match the same total compute FLOPs.
The optimizer is Adam (Kingma & Ba, 2014) with learning rate $5\times 10^{-4}$ and weight decay $5\times 10^{-5}$, using an inverse square root decay schedule with 5K warm-up and cool-down steps. This learning rate was chosen by sweeping across the values $\mathrm{lr}\in\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5}\}$ with $\mathrm{wd}=\mathrm{lr}/10$. The reported value yielded the best performance for the non-recursive baseline. The full training configuration along with a list of all the architectures we sweep across (i.e. signature and degree) can be found in Appendix A. Each model is trained on $16\times 16$ TPUv5 chips for approximately 2.2K core-hours. Models that failed OOM were excluded. Overall, we train 59 models of two sizes: 300M and 600M parameters. In all models, the embedding dimension is 1,536 and the MLP dimension is 6,144. Later, we present results using models with 1 billion parameters.
@@ Results.
One observation in Figure 2 is that the performance gap seems to widen in favor of RINS as more training compute is allocated. To verify this, we resume training the 600M-parameter models on 500B training tokens. Here, we only compare RINS with the non-recursive baseline and RAO, since RAO was identified in prior works to outperform other parameter-sharing strategies (Liu et al., 2024). Figure 4 shows that the performance gap continues to increase in favor of RINS. This improvement also persists downstream, as shown in Table 1.

### v3

@@ 3 Experimental Results
In this section, we study the impact of various parameter-sharing strategies in language modeling, following the taxonomy introduced in Section 2. We show how RINS emerges as a clear winner. All experiments are carried out using the Big Vision codebase (Beyer et al.,, 2022).
BL RAO RINS BL RAO RINS
OpenBookQA $37.1\pm 0.8$ $\bf 40.0\pm 0.6$ $39.0\pm 0.3$ BoolQ $53.1\pm 4.7$ $58.3\pm 0.6$ $\bf 59.5\pm 0.8$
PIQA $65.8\pm 3.4$ $68.8\pm 0.5$ $\bf 69.8\pm 0.3$ SIQA $40.2\pm 0.5$ $40.0\pm 0.5$ $\bf 40.9\pm 0.4$
HellaSwag $46.0\pm 1.0$ $49.6\pm 0.3$ $\bf 50.2\pm 0.5$ CommonSenseQA $29.0\pm 1.6$ $31.3\pm 1.7$ $\bf 32.4\pm 1.0$
Table 1: 0-shot evaluation in downstream common sense reasoning tasks. All models are 600M parameters, pretrained on the compute-equivalent of 500B tokens in the baseline (BL). The best signatures for RAO and RINS in Figure 3(a) are used in this evaluation. See Appendix C for details.
@@
First, we train a decoder-only transformer language model (Vaswani et al.,, 2017) with relative position embeddings and sequence packing. We use C4/English tokenizer with a vocabulary size of 32K. The model is trained on a mixture of C4 (Raffel et al.,, 2020) and SlimPajama (Soboleva et al.,, 2023) with equal weight using a batch size 1,024 and context length 1,024. The non-recursive baseline is trained for 200K steps, which amounts to about 200B training tokens. Other recursive models are trained on fewer tokens in order to match the same total compute FLOPs.
The optimizer is Adam (Kingma and Ba,, 2014) with learning rate $5\times 10^{-4}$ and weight decay $5\times 10^{-5}$, using an inverse square root decay schedule with 5K warm-up and 5K cool-down steps. This learning rate was chosen by sweeping across the values $\mathrm{lr}\in\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5}\}$ with $\mathrm{wd}=\mathrm{lr}/10$. The reported value yielded the best performance for the non-recursive baseline. The full training configuration along with a list of all the architectures we sweep across (i.e. signature and degree) can be found in Appendix E. Each model is trained on $16\times 16$ TPUv5 chips for approximately 2.2K core-hours. Models that failed OOM were excluded. Overall, we train 59 models of two sizes: 300M and 600M parameters, which include RAO (Liu et al.,, 2024) and latent recurrent thinking (Geiping et al.,, 2025) whose signature is ABrC. In all models, the embedding dimension is 1,536 and the MLP dimension is 6,144. In Figure 3(a), we present results using models with 1 billion parameters.
@@ Results.
One observation in Figure 2 is that the performance gap seems to widen in favor of RINS as more training compute is allocated. To verify this, we resume training the 600M-parameter models on 500B training tokens. Here, we only compare RINS with the non-recursive baseline and RAO, since RAO was identified in prior works to outperform other parameter-sharing strategies (Liu et al.,, 2024). Figure 3(a) shows that the performance gap continues to increase in favor of RINS. This improvement also persists downstream, as shown in Table 1.